increase of $25,000 after five years. He can also invest his money in risky assets where there are three outcomes: he has a 31 percent chance of losing $30,000, a 12 percent chance of gaining S15,000 and a 57 percent chance of gaining an unknown amount of money. What must the unknown amount of money equal if the expected value is to equal the guaranteed $25,000? Give your answer to the nearest whole dollar (with no decimal points, spaces, S signs, or commas in your answer)
Answer #1

Ans: 58018

Explanation:

Let the unknown amount is X

Expected value = -0.31(30,000) + 0.12(15,000) + 0.57(X)

25,000 = -9,300 + 1,800 + 0.57X

0.57X = 25,000 + 9,300 - 1,800 = 32,500

X = 32,500 / 0.57 = 57018

Thus, the unknown amount is 57018